OpenStudy (anonymous):

The domain of the following relation: R: {(3, −3), (1, 6), (−5, −4), (−10, 6)} is

OpenStudy (anonymous):

R:\[\left\{ \left( x1,y1 \right),\left( x2,y2 \right),\left( x3,y3 \right),\left( x4,y4 \right) etc \right\} is \left\{ x1,x2,x3,x4,etc \right\}\]

OpenStudy (anonymous):

IS that the answwer ? @surjithayer

OpenStudy (anonymous):

yes ,you are correct.

OpenStudy (anonymous):

wait what is the answer ? @surjithayer @jdoe0001

OpenStudy (jdoe0001):

well, check above :), the picture, see which is the domain in the set?

OpenStudy (jdoe0001):

$$ \text{The domain of the following relation: }\\ \\ R: {(\color{red}{3}, −3), (\color{red}{1}, 6), (\color{red}{−5}, −4), (\color{red}{−10}, 6)} $$ see it now?

OpenStudy (anonymous):

If f(x) = −4x + 1, find f(−5). @jdoe0001

OpenStudy (jdoe0001):

$$ f(\color{red}{x}) = −4\color{red}{x} + 1\\ f(\color{red}{-5}) = −4(\color{red}{-5}) + 1 $$

OpenStudy (jdoe0001):

they just mean, make x = -5